TWO ATHLETES PETER AND JOHN HAVE THE PROBABILITY OF 1/3 AND 3/4 RESPECTIVELY TO QUALIFY FOR THE FINALS OF A HIGH JUMP.IF THEIR ATTEMPTS ARE INDEPENDENT,DETERMINE THE PROBABILITY THAT,
BOTH WILL QUALIFY FOR THE FUNDS
1
Expert's answer
2012-11-12T12:28:09-0500
& P(BOTH WILL QUALIFY FOR THE FUNDS) =

= P(PETER WILL QUALIFY FOR THE FUNDS) * P(JOHN WILL QUALIFY FOR THE FUNDS) =

= 1/3 * 3/4 = 3/12.